OpenStudy (anonymous):

How do you algebraically determine the derivative of y=(sqrt(x))(x+1)?

OpenStudy (anonymous):

I got y' = (1/2)x^(-1/2), but the answer is y' = (3/2)x^(1/2) + (1/2)x^(-1/2)

OpenStudy (jusaquikie):

\[y = \frac{ \sqrt{x} }{ x+1 }\] well without using calculus you would need the rise over run which would be \[y' = \frac{ f(b) - f(a) }{b-a }\]

OpenStudy (jusaquikie):

for b and a you can use any numbers in the domain

OpenStudy (jusaquikie):

you may have see the formula like this before : \[m=\frac{ y _{2} -y_{1}}{ x_{2}-x_{1} }\]

OpenStudy (anonymous):

I think I got it, thanks guys.

OpenStudy (anonymous):

What do you mean by 'algebraically determine?'

OpenStudy (anonymous):

Meaning anything but graphically

OpenStudy (anonymous):

If you want derivative, I'd recommend product rule. This is differential calculus, yes?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

For \[\large y=(\sqrt{x})(x+1) \] is the function?

OpenStudy (anonymous):

Product rule: for y=u(x)*v(x) y' = u(x)*v'(x) + v(x)*u'(x)